OpenStudy (anonymous):

Solve each system by graphing help! 12. Y=1/3x+1 Y=-3x+11

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Hmm so it's pretty much the same type of problem as we did before :O Do you know how to graph each of these lines?

OpenStudy (anonymous):

we are still using y=mx+b and uhh brain fart

OpenStudy (anonymous):

Im sorry i just have a ton of work to do

zepdrix (zepdrix):

heh

zepdrix (zepdrix):

The lines will cross the y-axis at the b values.

OpenStudy (anonymous):

Got it! and the b value is 1 for the first one correct?

zepdrix (zepdrix):

Yah that sounds right :)

OpenStudy (anonymous):

YAY! I finally get something lol

zepdrix (zepdrix):

Hmm so we need another point on this line.\[\large y=\frac{1}{3}x+1\]What would be a good x value that we can plug in? We want to plug in a value that will deal with the fraction nicely.

OpenStudy (anonymous):

hmm

zepdrix (zepdrix):

We have a 1/3.. hmmm What happens to the 1/3 if we plug in x=3?

OpenStudy (anonymous):

OMGGG ! i was trying it to make sure it fit. sounds good.. we'll get 2

zepdrix (zepdrix):

Ok good! That gives us another point on our line. Now we can draw the line through these two points.